# Weak Convexity of Fisher Information Matrix and Superresolved Localization of Blinking Sources of Light

### v2

@@ II.3 Quantum Cramér-Rao bound
For a given state of the object, the conditional probabilities, Eq. (2), and therefore the parameter variances calculated from them, depend on four quantities: $\bm{\rho}_{P}$, $\bm{\rho}_{E}$, $\bm{U}_{\theta}$, and $\bm{\Lambda}(\bm{r})$. Each of these quantities can be manipulated to optimize the estimation. In particular, optimization of $\bm{\Lambda}(\bm{r})$ corresponds to the choice of the best measurement for the given state of the probe $\tilde{\bm{\rho}}_{P}(\theta)$. For this purpose, quantum Fisher information matrix $\mathbf{Q}(\theta)$ is defined as [39, 40, 41]
$\mathbf{Q}_{ij}(\theta)=\frac{1}{2}\operatorname{Tr}\left\{\left(\mathcal{L}_{i}\mathcal{L}_{j}+\mathcal{L}_{j}\mathcal{L}_{i}\right)\tilde{\bm{\rho}}_{P}(\theta)\right\},$ (14)
@@
$\frac{\partial}{\partial\theta_{i}}\tilde{\bm{\rho}}_{P}(\theta)=\frac{1}{2}\left[\mathcal{L}_{i}\tilde{\bm{\rho}}_{P}(\theta)+\tilde{\bm{\rho}}_{P}(\theta)\mathcal{L}_{i}\right].$ (15)
Quantum Fisher information matrix satisfies the condition $\mathbf{F}^{-1}(\theta)\geq\mathbf{Q}^{-1}(\theta)$, where $\mathbf{F}^{-1}(\theta)$ is calculated for the same state of the probe with an arbitrary POVM $\bm{\Lambda}(\mathbf{r})$. Moreover, there is often an optimal POVM $\bm{\Lambda}_{\text{opt}}(\mathbf{r})$ (but not always [25]) for which $\mathbf{F}_{\text{opt}}^{-1}(\theta)=\mathbf{Q}^{-1}(\theta)$. This means that the quantity $\operatorname{Tr}\{\mathbf{Q}^{-1}(\theta)\}/M$ gives the lower bound of the average variance for the optimized measurement. As mentioned above, manipulation of $\bm{\rho}_{P}$ by entangling the probes is known to further minimize the estimation error in some tasks of quantum metrology [23]. In this paper, we show that the superresolution obtained by making the sources blink corresponds to manipulation of $\bm{\rho}_{O}(\theta)$ without affecting the value of the parameter $\theta$.
@@ III Additivity and weak convexity of Fisher information matrix
Given two positive integers $\ell$ and $M$ and a random variable $X\in\mathbb{R}^{\ell}$ depending on parameter $\theta\in\mathbb{R}^{M}$, so that the probability density for $X$ under condition that the parameter is $\theta$ is $q(X|\theta)$, the Fisher information matrix for the information contained in $X$ about $\theta$ is the $M\times M$-matrix-valued functional $\mathcal{F}_{M}^{(\ell)}\left[q(X|\theta)\right]$, whose matrix elements are
@@
$\mathcal{F}_{M}^{(\ell_{1}+\ell_{2})}\left[q_{1}(X_{1}|\theta)q_{2}(X_{2}|\theta)\right]$ (17)
$=\mathcal{F}_{M}^{(\ell_{1})}\left[q_{1}(X_{1}|\theta)\right]+\mathcal{F}_{M}^{(\ell_{2})}\left[q_{2}(X_{2}|\theta)\right],$
@@
for any $10<\gamma<$ 1 and $q_{1}(X|\theta)\not\equiv q_{2}(X|\theta)$.
@@
Given two positive integers $K\geq 2$ and $d$, in the general case of $M$ parameters split into $K$ groups of $d$ parameters each, $\theta=\{\bar{\theta}_{1},...,\bar{\theta}_{K}\}$, where $\bar{\theta}_{k}\in\mathbb{R}^{d}$, $M=Kd$, and given real numbers $\mu_{1}$,…,$\mu_{K}$ belonging to $[0,1)$ and summing up to unity, the Fisher information matrix possesses the following properties.
@@
    $\mathcal{F}_{M}^{(\ell)}\left[\frac{1}{K}\sum\limits_{k=1}^{K}q_{k}(X|\theta)\right]<\frac{1}{K}\sum\limits_{k=1}^{K}\mathcal{F}_{M}^{(\ell)}\left[q_{k}(X|\theta)\right],$ (19)
@@
    $\operatorname{Tr}\left\{\mathcal{F}_{M}^{(\ell)}\left[\sum\limits_{k=1}^{K}\mu_{k}q_{k}(X|\theta)\right]\right\}<\sum\limits_{k=1}^{K}\mu_{k}\operatorname{Tr}\left\{\mathcal{F}_{M}^{(\ell)}\left[q_{k}(X|\theta)\right]\right\}.$ (21)
@@
Let us denote the matrix in the left-hand side of Eq. (21) by $F$ and that in the right-hand side by $\mathring{F}=\sum_{k}\mu_{k}\mathcal{F}_{M}^{(\ell)}\left[q_{k}(X|\theta)\right]$. $F$ is a real symmetric matrix and as such can be brought to a diagonal form by an orthogonal rotation in the parameter space: $F=OD_{F}O^{T}$, where $D_{F}$ is diagonal and $O$ is orthogonal. Consider a linear transformation from the (column) vector of parameters $\theta=\{\theta_{1},...,\theta_{M}\}$ to the new vector of parameters $\varphi=\{\varphi_{1},...,\varphi_{M}\}$, $\varphi=O^{T}\theta$. The conditional probability density of $X$ for given $\varphi$ is $\tilde{q}(X|\varphi)=q(X|O\varphi)$. Let us denote the Fisher information matrix for the information contained in $X$ about $\varphi$ by $\tilde{F}$. According to Definition 4, its elements are
@@
or $\tilde{F}=O^{T}FO=D_{F}$. This means that $\varphi$ can be considered as a vector of “eigenparameters” of the matrix $F$. Let us fix all eigenparameters except for one, $\varphi_{i}$, and write the conditional probability density of $X$ for given $\varphi_{i}$ as $\tilde{q}^{\prime}(X|\varphi_{i})=\tilde{q}(X|\varphi)$. The Fisher information contained in $X$ about $\varphi_{i}$ is
@@
Applying Eq. (35) to each side of this inequality, we obtain
@@
$\left\{O^{T}\left(\mathring{F}-F\right)O\right\}_{ii}>0.$ (26)
In the case of part A, we have $\mathring{F}=f_{0}I_{M}/K$, where $I_{M}$ is the unit $M\times M$ matrix, which can be seen applying the condition of equal weights, $\mu_{k}=1/K$, and then the identity $\sum_{k}\mathcal{F}_{M}^{(\ell)}\left[q_{k}(X|\theta)\right]_{mn}=f_{0}\delta_{mn}$ following from Eq. (20). Since inequality (26) holds for any $i$ and both matrices $O^{T}\mathring{F}O=\mathring{F}$ and $O^{T}FO=D_{F}$ are diagonal, this inequality can be rewritten as $O^{T}\left(\mathring{F}-F\right)O>0$ or $F<\mathring{F}$, which gives Eq. (36). Part A of the theorem is proved.
In the general case, we sum Eq. (26) over $i$ and use the invariance of the trace with respect to orthogonal rotations to obtain $\operatorname{Tr}F<\operatorname{Tr}\mathring{F}$, which coincides with Eq. (21) and proves part B. ∎
Note that the area of applicability of the matrix convexity is slightly wider than that stated in part A of Theorem 1. Indeed, what we need for the proof, is that $\mathring{F}$ is proportional to the unit matrix, which can be reached by replacing $f_{0}$ by $f_{k}$ in Eq. (20) and demanding, instead of equal weights, that $\mu_{k}f_{k}$ is independent of $k$.
@@ IV Localization of blinking sources
$\bm{\rho}_{O}^{\text{coflu}}=\left(\sum_{k}\mu_{k}\rho_{Ok}\right)^{\otimes N}.$ (27)
All other quantities in the right-hand side of Eq. (2) correspond to independent identical generalized measurements (IIGM), that is they are simple $N$th powers of the single-sample quantities:
$\bm{\rho}_{P}=\rho_{P}^{\otimes N},\,\,\bm{\rho}_{E}=\rho_{E}^{\otimes N},\,\,\bm{U}=U^{\otimes N},\,\,\bm{\Lambda}(\bm{r})=\prod\limits_{n}\Lambda(r_{n}).$ (28)
Substituting these expressions into Eq. (2), we obtain $\bm{p}(\bm{r}|\theta)=\prod_{n}p(r_{n}|\theta)$, where
@@
is the single-sample conditional probability of measurement outcome $r=\{x,y\}$, which is the location of the detected photon. Substituting this probability into Eq. (4), we find the Fisher information matrix in the form $\mathbf{F}_{\text{coflu}}(\theta)=NF_{\text{coflu}}(\theta)$, where $F_{\text{coflu}}(\theta)$ is the single-sample Fisher information matrix with the elements
[…]
@@
$\bm{\rho}_{O}^{\text{blink}}=\rho_{O1}^{\otimes N_{1}}\otimes...\otimes\rho_{OK}^{\otimes N_{K}}.$ (31)
All other quantities in the right-hand side of Eq. (2) are simple $N$th powers of the single-sample quantities, as in the previous scenario. Substituting these expressions into Eq. (2), we obtain $\bm{p}(\bm{r}|\theta)=\prod_{k=1}^{K}\bm{p}_{k}(\bm{r}_{k}|\theta)$, where $\bm{r}_{k}$ is a part of $\bm{r}$ with the outcomes in the $k$th super-window and $\bm{p}_{k}(\bm{r}_{k}|\theta)=\prod_{l=1}^{N_{k}}p_{k}(r_{kl}|\theta)$ with
@@
the single-sample conditional probability of measurement outcome when the $k$th source is only active. Here, the index $kl$ means the $l$th element in the $k$th super-window and equals $kl=l+\sum_{m=1}^{k-1}N_{m}$. Substituting this probability into Eq. (4), we find the Fisher information matrix in the form $\mathbf{F}_{\text{blink}}(\theta)=\sum_{k}N_{k}F^{[k]}(\theta)$, where $F^{[k]}(\theta)$ is the single-sample Fisher information matrix in the $k$th window, where just the $k$th source is active:
[…]
Actually, $F^{[k]}(\theta)$ depends on the location $\bar{\theta}_{k}=\bar{r}_{k}$ of the $k$th source only, because other sources are inactive in this super-window.
@@
$p(Rr|R\bar{r})$ $=$ $p(r|\bar{r}).$ (35)
where $\bar{r}_{0}=\{\bar{x}_{0},\bar{y}_{0}\}$ is an arbitrary real vector and $R$ is a rotation matrix on the object plane.
@@
In translation and rotation invariant microscopy, the Fisher information matrix on the position of a single source $\bar{r}=\{\bar{x},\bar{y}\}$ is proportional to the unit $2\times 2$ matrix $I$,
$\mathcal{F}_{2}^{(2)}\left[p(r|\bar{r})\right]=f_{0}I,$ (36)
@@
Rotating the parameters as $\bar{r}\to R\bar{r}$, we find from Eq. (16) the corresponding rotation of the Fisher information matrix, $\mathcal{F}_{2}^{(2)}\left[p(r|R\bar{r})\right]=R\mathcal{F}_{2}^{(2)}\left[p(r|\bar{r})\right]R^{T}$. Since the Jacobian corresponding to a rotation of integration coordinates is $|\det R|=1$, we also have $\mathcal{F}_{2}^{(2)}\left[p(Rr|R\bar{r})\right]=\mathcal{F}_{2}^{(2)}\left[p(r|R\bar{r})\right]$. Combining these two equalities with Eq. (35), we obtain invariance of $\mathcal{F}_{2}^{(2)}\left[p(r|\bar{r})\right]$ to rotations, which means that it is proportional to the unit matrix $I$. Substituting Eq. (34) into Eq. (16), we find $\mathcal{F}_{2}^{(2)}\left[p(r|\bar{r}+\bar{r}_{0})\right]=\mathcal{F}_{2}^{(2)}\left[p(r|\bar{r})\right]$, which means that $f_{0}$ is independent of $\bar{r}$. ∎
@@
For $\mathbf{F}_{\text{coflu}}(\theta)$ defined by Eqs. (2) and (4) with the object state (27) and $\mathbf{F}_{\text{blink}}(\theta)$ defined by Eqs. (2) and (4) with the object state (31) and IIGM defined by Eq. (28), the following inequalities hold for any $\rho_{Ok}$, $\rho_{P}$, $\rho_{E}$, $U$, and $\Lambda(r)$.
@@
    $\mathbf{F}_{\text{blink}}(\theta)>\mathbf{F}_{\text{coflu}}(\theta).$ (37)
@@
    $\operatorname{Tr}\{\mathbf{F}_{\text{blink}}(\theta)\}>\operatorname{Tr}\{\mathbf{F}_{\text{coflu}}(\theta)\}.$ (38)
@@
Substituting Eqs. (27) and (28) into Eq. (2) and using the property $\operatorname{Tr}\{A\otimes B\}=\operatorname{Tr}A\operatorname{Tr}B$, we obtain $\bm{p}(\bm{r}|\theta)=\prod_{n}p(r_{n}|\theta)$ with $p(r|\theta)$ defined by Eq. (29), which can be rewritten as
@@
where $p_{k}(r|\theta)$ is defined by Eq. (32). Substituting this form of $\bm{p}(\bm{r}|\theta)$ into Eq. (4) and using the additivity of Fisher information by Lemma 1, we obtain
$\mathbf{F}_{\text{coflu}}(\theta)=NF_{\text{coflu}}(\theta),$ (40)
where $F_{\text{coflu}}(\theta)$ is a single-sample Fisher information matrix calculated as a functional of $p(r|\theta)$, $F_{\text{coflu}}(\theta)=\mathcal{F}_{M}^{(2)}[p(r|\theta)]$, where the upper index $\ell=2$ is the dimension of the vector $r=\{x,y\}$.
On the other hand, substituting Eqs. (31) and (28) into Eq. (2), we obtain $\bm{p}(\bm{r}|\theta)=\prod_{k=1}^{K}\bm{p}_{k}(\bm{r}_{k}|\theta)$, where $\bm{p}_{k}(\bm{r}_{k}|\theta)=\prod_{l=1}^{N_{k}}p_{k}(r_{kl}|\theta)$ and $p_{k}(r_{kl}|\theta)$ is given by Eq. (32). Substituting this form of $\bm{p}(\bm{r}|\theta)$ into Eq. (4) and using the additivity of Fisher information by Lemma 1, we obtain
$\mathbf{F}_{\text{blink}}(\theta)=\sum_{k}N_{k}F^{[k]}(\theta),$ (41)
where $F^{[k]}(\theta)$ is given by Eq. (33) and calculated as a functional of $p_{k}(r|\theta)$, $F^{[k]}(\theta)=\mathcal{F}_{M}^{(2)}[p_{k}(r|\theta)]$. Finally, we apply the equality $N_{k}=N\mu_{k}$. In the case of $N_{k}=N/K$, using the convexity of the Fisher information matrix by part A of Theorem 1, we obtain Eq. (37). The condition (20) is satisfied according to Lemma 3. In the case of arbitrary $N_{k}$, using the convexity of the trace of the Fisher information matrix by part B of Theorem 1, we obtain Eq. (38). ∎
If the matrices $A$ and $B$ are invertible and positive, $A>B$ is equivalent to $A^{-1}<B^{-1}$ [35]. As a corollary to Theorem 2, we have $\mathbf{F}_{\text{blink}}^{-1}(\theta)<\mathbf{F}_{\text{coflu}}^{-1}(\theta)$, and therefore, for a saturated Cramér-Rao bound, in the case of equal brightnesses,
$H_{\text{blink}}(\hat{\theta})>H_{\text{coflu}}(\hat{\theta}),$ (42)
where $H(\hat{\theta})$ may be any of the three measures of localization efficiency introduced in Sec. II.2, $H_{\text{tot}}(\hat{\theta})$, $H_{\text{ind}}(\hat{\theta})$, or $H_{\text{eig}}(\hat{\theta})$. Moreover, in the case of arbitrary brightnesses, Eq. (42) also holds for $H_{\text{eig}}(\hat{\theta})$.

### v3

@@ II.3 Quantum Cramér-Rao bound
For a given state of the object, the conditional probabilities, Eq. (2), and therefore the parameter variances calculated from them, depend on four quantities: $\underline{\rho}_{P}$, $\underline{\rho}_{E}$, $\underline{U}_{\theta}$, and $\underline{\Lambda}(\underline{r})$. Each of these quantities can be manipulated to optimize the estimation. In particular, optimization of $\underline{\Lambda}(\underline{r})$ corresponds to the choice of the best measurement for the given state of the probe $\tilde{\underline{\rho}}_{P}(\theta)$. For this purpose, quantum Fisher information matrix $\underline{\mathbf{Q}}(\theta)$ is defined as [39, 40, 41]
$\underline{Q}_{ij}(\theta)=\frac{1}{2}\operatorname{Tr}\left\{\left(\mathcal{L}_{i}\mathcal{L}_{j}+\mathcal{L}_{j}\mathcal{L}_{i}\right)\tilde{\underline{\rho}}_{P}(\theta)\right\},$ (14)
@@
$\frac{\partial}{\partial\theta_{i}}\tilde{\underline{\rho}}_{P}(\theta)=\frac{1}{2}\left[\mathcal{L}_{i}\tilde{\underline{\rho}}_{P}(\theta)+\tilde{\underline{\rho}}_{P}(\theta)\mathcal{L}_{i}\right].$ (15)
Quantum Fisher information matrix satisfies the condition $\underline{\mathbf{F}}^{-1}(\theta)\geq\underline{\mathbf{Q}}^{-1}(\theta)$, where $\underline{\mathbf{F}}^{-1}(\theta)$ is calculated for the same state of the probe with an arbitrary POVM $\underline{\Lambda}(\underline{r})$. Moreover, there is often an optimal POVM $\underline{\Lambda}_{\text{opt}}(\underline{r})$ (but not always [25]) for which $\underline{\mathbf{F}}_{\text{opt}}^{-1}(\theta)=\underline{\mathbf{Q}}^{-1}(\theta)$. This means that the quantity $\operatorname{Tr}\{\underline{\mathbf{Q}}^{-1}(\theta)\}/M$ gives the lower bound of the average variance for the optimized measurement. As mentioned above, manipulation of $\underline{\rho}_{P}$ by entangling the probes is known to further minimize the estimation error in some tasks of quantum metrology [23]. In this paper, we show that the superresolution obtained by making the sources blink corresponds to manipulation of $\underline{\rho}_{O}(\theta)$ without affecting the value of the parameter $\theta$.
@@ III Additivity and weak convexity of Fisher information matrix
[…]
@@
[…]
@@
for any $0<\gamma<1$ and $q_{1}(X|\theta)\not\equiv q_{2}(X|\theta)$.
@@
Given two positive integers $K\geq 2$ and $d$, in the general case of $M$ parameters split into $K$ groups of $d$ parameters each, $\theta=\{\bar{\theta}_{1},...,\bar{\theta}_{K}\}$, where $\bar{\theta}_{k}\in\mathbb{R}^{d}$, $M=Kd$, and given real numbers $\mu_{1}$$\mu_{K}$ belonging to $[0,1)$ and summing up to unity, the Fisher information matrix possesses the following properties.
@@
[…]
@@
[…]
@@
Let us denote the matrix in the left-hand side of Eq. (21) by $\mathbf{F}$ and that in the right-hand side by $\mathring{\mathbf{F}}=\sum_{k}\mu_{k}\bm{\mathcal{F}}_{M}^{(\ell)}\left[q_{k}(X|\theta)\right]. 𝐅\mathbf${F} is a real symmetric matrix and as such can be brought to a diagonal form by an orthogonal rotation in the parameter space: $\mathbf{F}=\mathbf{O}\mathbf{D}_{F}\mathbf{O}^{T}$, where $\mathbf{D}_{F}$ is diagonal and $\mathbf{O}$ is orthogonal. Consider a linear transformation from the (column) vector of parameters $\theta=\{\theta_{1},...,\theta_{M}\}$ to the new vector of parameters $\varphi=\{\varphi_{1},...,\varphi_{M}\}$, $\varphi=\mathbf{O}^{T}\theta$. The conditional probability density of $X$ for given $\varphi$ is $\tilde{q}(X|\varphi)=q(X|\mathbf{O}\varphi)$. Let us denote the Fisher information matrix for the information contained in $X$ about $\varphi$ by $\tilde{\mathbf{F}}$. According to Definition 4, its elements are
@@
or $\tilde{\mathbf{F}}=\mathbf{O}^{T}\mathbf{F}\mathbf{O}=\mathbf{D}_{F}$. This means that $\varphi$ can be considered as a vector of “eigenparameters” of the matrix $\mathbf{F}$. Let us fix all eigenparameters except for one, $\varphi_{i}$, and write the conditional probability density of $X$ for given $\varphi_{i}$ as $\tilde{q}^{\prime}(X|\varphi_{i})=\tilde{q}(X|\varphi)$. The Fisher information contained in $X$ about $\varphi_{i}$ is
@@
Applying Eq. (III) to each side of this inequality, we obtain
@@
$\left\{\mathbf{O}^{T}\left(\mathring{\mathbf{F}}-\mathbf{F}\right)\mathbf{O}\right\}_{ii}>0.$ (26)
In the case of part A, we have $\mathring{\mathbf{F}}=f_{0}\mathbf{I}_{M}/K$, where $\mathbf{I}_{M}$ is the unit $M\times M$ matrix, which can be seen applying the condition of equal weights, $\mu_{k}=1/K$, and then the identity $\sum_{k}\mathcal{F}_{M}^{(\ell)}\left[q_{k}(X|\theta)\right]_{mn}=f_{0}\delta_{mn}$ following from Eq. (20). Since inequality (26) holds for any $i$ and both matrices $\mathbf{O}^{T}\mathring{\mathbf{F}}\mathbf{O}=\mathring{\mathbf{F}}$ and $\mathbf{O}^{T}\mathbf{F}\mathbf{O}=\mathbf{D}_{F}$ are diagonal, this inequality can be rewritten as $\mathbf{O}^{T}\left(\mathring{\mathbf{F}}-\mathbf{F}\right)\mathbf{O}>0$ or $\mathbf{F}<\mathring{\mathbf{F}}$, which gives Eq. (19). Part A of the theorem is proved.
In the general case, we sum Eq. (26) over $i$ and use the invariance of the trace with respect to orthogonal rotations to obtain $\operatorname{Tr}\mathbf{F}<\operatorname{Tr}\mathring{\mathbf{F}}$, which coincides with Eq. (21) and proves part B. ∎
Note that the area of applicability of the matrix convexity is slightly wider than that stated in part A of Theorem 1. Indeed, what we need for the proof, is that $\mathring{\mathbf{F}}$ is proportional to the unit matrix, which can be reached by replacing $f_{0}$ by $f_{k}$ in Eq. (20) and demanding, instead of equal weights, that $\mu_{k}f_{k}$ are independent of $k$.
@@ IV Localization of blinking sources
$\underline{\rho}_{O}^{\text{coflu}}=\left(\sum_{k}\mu_{k}\rho_{Ok}\right)^{\otimes N}.$ (27)
All other quantities in the right-hand side of Eq. (2) correspond to independent identical generalized measurements, that is they are simple $N$th powers of the single-sample quantities:
$\underline{\rho}_{P}=\rho_{P}^{\otimes N},\,\,\underline{\rho}_{E}=\rho_{E}^{\otimes N},\,\,\underline{U}=U^{\otimes N},\,\,\underline{\Lambda}(\underline{r})=\prod\limits_{n}\Lambda(r_{n}).$ (28)
Substituting these expressions into Eq. (2), we obtain $\underline{p}(\underline{r}|\theta)=\prod_{n}p(r_{n}|\theta)$, where
@@
is the single-sample conditional probability of measurement outcome $r=\{x,y\}$, which is the location of the detected photon. Substituting this probability into Eq. (4), we find the Fisher information matrix in the form $\underline{\mathbf{F}}_{\text{coflu}}(\theta)=N\mathbf{F}_{\text{coflu}}(\theta)$, where $\mathbf{F}_{\text{coflu}}(\theta)$ is the single-sample Fisher information matrix with the elements
$F_{\text{coflu},ij}(\theta)=\int\frac{1}{p(r|\theta)}\frac{\partial p(r|\theta)}{\partial\theta_{i}}\frac{\partial p(r|\theta)}{\partial\theta_{j}}dr.$ (30)
@@
$\underline{\rho}_{O}^{\text{blink}}=\rho_{O1}^{\otimes N_{1}}\otimes...\otimes\rho_{OK}^{\otimes N_{K}}.$ (31)
All other quantities in the right-hand side of Eq. (2) are simple $N$th powers of the single-sample quantities, as in the previous scenario. Substituting these expressions into Eq. (2), we obtain $\underline{p}(\underline{r}|\theta)=\prod_{k=1}^{K}\underline{p}_{k}(\underline{r}_{k}|\theta)$, where $\underline{r}_{k}$ is a part of $\underline{r}$ with the outcomes in the $k$th super-window and $\underline{p}_{k}(\underline{r}_{k}|\theta)=\prod_{l=1}^{N_{k}}p_{k}(r_{kl}|\theta)$ with
@@
the single-sample conditional probability of measurement outcome when the $k$th source is only active. Here, the index $kl$ means the $l$th element in the $k$th super-window and equals $kl=l+\sum_{m=1}^{k-1}N_{m}$. Substituting this probability into Eq. (4), we find the Fisher information matrix in the form $\underline{\mathbf{F}}_{\text{blink}}(\theta)=\sum_{k}N_{k}\mathbf{F}^{[k]}(\theta)$, where $\mathbf{F}^{[k]}(\theta)$ is the single-sample Fisher information matrix in the $k$th window, where just the $k$th source is active:
$F_{ij}^{[k]}(\theta)=\int\frac{1}{p_{k}(r|\theta)}\frac{\partial p_{k}(r|\theta)}{\partial\theta_{i}}\frac{\partial p_{k}(r|\theta)}{\partial\theta_{j}}dr.$ (33)
Actually, $\mathbf{F}^{[k]}(\theta)$ depends on the location $\bar{\theta}_{k}=\bar{r}_{k}$ of the $k$th source only, because other sources are inactive in this super-window.
@@
$p(\mathbf{R}r|\mathbf{R}\bar{r})$ $=$ $p(r|\bar{r}).$ (35)
where $\bar{r}_{0}=\{\bar{x}_{0},\bar{y}_{0}\}$ is an arbitrary real vector and $\mathbf{R}$ is a rotation matrix on the object plane.
@@
In translation and rotation invariant microscopy, the Fisher information matrix on the position of a single source $\bar{r}=\{\bar{x},\bar{y}\}$ is proportional to the unit $2\times 2$ matrix $\mathbf{I}$,
[…]
@@
[…]
@@
For $\underline{\mathbf{F}}_{\text{coflu}}(\theta)$ defined with the object state (27), $\underline{\mathbf{F}}_{\text{blink}}(\theta)$ defined with the object state (31), and all generalized measurements being independent and identical, the following inequalities hold for any $\rho_{Ok}$, $\rho_{P}$, $\rho_{E}$, $U$, and $\Lambda(r)$.
@@
    $\underline{\mathbf{F}}_{\text{blink}}(\theta)>\underline{\mathbf{F}}_{\text{coflu}}(\theta).$ (37)
@@
    $\operatorname{Tr}\{\underline{\mathbf{F}}_{\text{blink}}(\theta)\}>\operatorname{Tr}\{\underline{\mathbf{F}}_{\text{coflu}}(\theta)\}.$ (38)
@@
Substituting Eqs. (27) and (28) into Eq. (2) and using the property $\operatorname{Tr}\{\mathbf{A}\otimes\mathbf{B}\}=\operatorname{Tr}\mathbf{A}\operatorname{Tr}\mathbf{B}$, we obtain $\underline{p}(\underline{r}|\theta)=\prod_{n}p(r_{n}|\theta)$ with $p(r|\theta)$ defined by Eq. (29), which can be rewritten as
@@
where $p_{k}(r|\theta)$ is defined by Eq. (32). Substituting this form of $\underline{p}(\underline{r}|\theta)$ into Eq. (4) and using the additivity of Fisher information by Lemma 1, we obtain
$\underline{\mathbf{F}}_{\text{coflu}}(\theta)=N\mathbf{F}_{\text{coflu}}(\theta),$ (40)
[…]
On the other hand, substituting Eqs. (31) and (28) into Eq. (2), we obtain $\underline{p}(\underline{r}|\theta)=\prod_{k=1}^{K}\underline{p}_{k}(\underline{r}_{k}|\theta)$, where $\underline{p}_{k}(\underline{r}_{k}|\theta)=\prod_{l=1}^{N_{k}}p_{k}(r_{kl}|\theta)$ and $p_{k}(r_{kl}|\theta)$ is given by Eq. (32). Substituting this form of $\underline{p}(\underline{r}|\theta)$ into Eq. (4) and using the additivity of Fisher information by Lemma 1, we obtain
$\underline{\mathbf{F}}_{\text{blink}}(\theta)=\sum_{k}N_{k}\mathbf{F}^{[k]}(\theta),$ (41)
where $\mathbf{F}^{[k]}(\theta)$ is given by Eq. (33) and calculated as a functional of $p_{k}(r|\theta)$, $\mathbf{F}^{[k]}(\theta)=\bm{\mathcal{F}}_{M}^{(2)}[p_{k}(r|\theta)]. F$inally, we apply the equality $N_{k}=N\mu_{k}$. In the case of $N_{k}=N/K$, using the convexity of the Fisher information matrix by part A of Theorem 1, we obtain Eq. (37). The condition (20) is satisfied according to Lemma 3. In the case of arbitrary $N_{k}$, using the convexity of the trace of the Fisher information matrix by part B of Theorem 1, we obtain Eq. (38). ∎
If the matrices $\mathbf{A}$ and $\mathbf{B}$ are invertible and positive, $\mathbf{A}>\mathbf{B}$ is equivalent to $\mathbf{A}^{-1}<\mathbf{B}^{-1}$ [35]. As a corollary to Theorem 2, we have $\underline{\mathbf{F}}_{\text{blink}}^{-1}(\theta)<\underline{\mathbf{F}}_{\text{coflu}}^{-1}(\theta)$, and therefore, for a saturated Cramér-Rao bound, in the case of equal brightnesses,
[…]